I think I used the wrong email so I need to try this again.
Simplify the exponential expression.
x^9y^0
Thanks again

Answer by stanbon(75887)

x^9y^0
--------------
a^0 = 1 for all numbers "a" except 0
---------------
Your Problem:
= x^9 * 1
= x^9
=======
Cheers,
Stan H.